OpenStudy (anonymous):

The density of the gaseous element is 5 times that of oxygen under similar conditions. if the molecule is triatomic . what will be its atomic mass?

OpenStudy (anonymous):

under similar conditions, density is 5 times and the molecule is triatomic. I'd say the atomic mass is 10/3 times of oxygen atomic mass

OpenStudy (anonymous):

The answe should be 53.33

OpenStudy (foolaroundmath):

PV = nRT => PV = mRT/M (m = mass of gas, M = molecular mass) => P = (m/V)RT/M => P = dRT/M (d = m/V = density) So, P/RT = d/M Under similar conditions, d/M remains constant For oxygen, d/M = d(oxygen)/32 (Oxygen is diatomic => molecular mass = 2*atomic mass = 2*16) For our element, d/M = 5d(oxygen)/(3*atomic mass) These two are equal => 5/(3*atomic mass) = 1/32 => atomic mass = 5*32/3 = 53.33 amu